Circular No. 8423

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2004fd IN NGC 1060 Further to IAUC 8417, T. Boles reports his discovery of an apparent

supernova (mag 15.7) on unfiltered CCD images taken on Oct. 21.835 and 24.840 UT. The new object is located at α = 2^{h}43^{m}15^{s}.25, δ = +32^{o}25′25′′.6, which is ≈ 2′′.5 east and 4′′.3 south of the center of NGC 1060. SN 2004fd is not present on Boles’ images from Aug. 25, Sept. 8, and Oct. 5 (limiting mag 19.5) or on a Quick-V survey plate from 1982 (limiting mag 19.0).

V838 MONOCEROTIS N. M. Ashok and D. P. K. Banerjee, Physical Research Laboratory,

Ahmedabad, report the following near-infared magnitudes of V838 Mon, obtained following conjunction with the sun, using the 1.2-m Mt. Abu telescope: Oct. 4.021 UT, J = 7.50, H = 6.19, K = 5.52 (typical errors 0.05 mag). For comparison, magnitudes on Mar. 9.646 were J = 7.14, H = 5.96, and K = 5.3. Spectra (range 1.08–2.35 μm) obtained on Oct. 4 are reasonably similar to spectra taken earlier in 2004. The (1,0) and (4,0) A-X bands of AlO are prominently seen along with the deep absorption in the first-overtone band of CO at 2.29 μm. In addition, several second-overtone bands of CO (starting from the 3-0 to the 10-7 transition) are also seen in the H band.

COMET C/2004 U2 (SOHO) Further to IAUC 8419, another SOHO comet found on C3 coronagraph

images, this time by H. Otterstedt (‘discovery’ observation below), which K. Battams has found to have peaked near mag 7 before rapidly fading over the two-day arc of observation. The astrometry (measured by Battams, reduced by B. G. Marsden), orbital elements by Marsden (T = Oct. 18.0 TT, q = 0.066 AU, i = 84^{o}), and an ephemeris appear on MPEC 2004-U26.

2004 UT  α_{2000}  δ_{2000}
Oct. 16.596  13^{h}51.^{m}5  − 6^{o}45′

COMET C/2004 Q2 (MACHHOLZ) Total-magnitude and coma-diameter estimates: Sept. 16.70 UT, 9.9,

2′ (S. T. Rae, Hamilton, New Zealand, 0.20-m reflector); Oct. 11.71, 8.8, 3′.5 (K. Yoshimoto, Yamaguchi, Japan, 20×100 binoculars); 22.82, 7.7, 7′ (A. Pearce, Noble Falls, W. Australia, 8×40 binoculars).

2004 October 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green